# A Review of Graph-Powered Data Quality Applications for IoT Monitoring Sensor Networks

### v1

@@ 3.3.1 GNNs: Building Modules
Figure 9: Example of a GCN architecture at the node level, where a graph signal $\mathbf{x}$ is reconstructed $\tilde{\mathbf{x}}$. The input are the graph signal and the graph shift matrix $\mathbf{S}$, and $\sigma(\cdot)$ denotes an activation function. This is a classic architecture composed of graph convolutional layers, it can be seen as a graph autoencoder. More generally, a GNN architecture can include pooling layers as well as readout layers and classic layers, e.g., fully-connected or convolutional layers.
@@
3.3.2 Convolutional Graph Neural Networks
@@
where $\mathbf{h}^{(k)}_{c_{k}}{\in}\mathbb{R}^{N}$ is the $c_{l}$ (channel) hidden state at layer $k$, $\sigma(\cdot)$ is the activation function, $c_{k}{\in}\mathbb{N}^{+}$ is the number of channels in layer $k$, $agg(\cdot)$ is an aggregation function used to aggregate the different filtered versions of graph signal $\mathbf{h}_{i}^{(k-1)}{\in}\mathbb{R}^{N}$ (e.g., sum or average), and $\boldsymbol{\theta}_{ij}^{(k)}{\in}\mathbb{R}^{K_{ij}}$ is the convolution function associated with a set of learnable parameters. Different types of networks and convolutions assume a different parametrization of the learnable parameters $\boldsymbol{\theta}$. For instance, the graph convolutional network (GCN) [96] and Chebnet [35] use the Chebyshev polynomials to approximate the filtering, the CayleyNet [103] uses the Cayley polynomials to approximate the filtering, the Graph Wavelet Neural Network (GWNN) [104] uses wavelets to approximate the convolution operation, and the graph diffusion neural network (GDNN) [105] uses a graph diffusion convolution to implement the message passing operation, combining strengths of spatial and spectral methods. For instance, the implementation of the spectral graph convolution using the truncated expansion of the Chebyshev polynomials results in [35]:
@@
where $f(\cdot)$ is a function that combines the value at node $n$ with a function of the vicinity, $agg_{\forall i}(\cdot)$ is a function that aggregates the information of neighboring nodes of $\mathcal{N}(n)$, and ||||| | denotes a concatenation operation. Function $f$ combines the aggregated spatial information at node $n$ ($\mathbf{a}^{(k-1)}(n)$) with the previous hidden state $\mathbf{h}^{k-1}(n)$ at node $n$. Both $agg$ and $f$ can contain a set of learnable parameters $\boldsymbol{\theta}_{i}^{agg}$ and $\boldsymbol{\theta}_{j}^{f}$. This framework provides great generalization capabilities allowing the definition of aggregation and combination functions according to the data. Graph attention networks (GAT) [109] make use of the node spatial notion and graph structure to define the graph-based analogs of attention mechanisms.
3.3.3 Graph Autoencoders
GAEs are widely used in applications such as outlier detection and missing value imputation. GAEs are trained to reconstruct the input signal, and in the process, a latent representation of the data is learned. To force the neural network to learn a latent representation, the dimensions of the hidden features of the encoder are decreased to produce a bottleneck effect, in what is called an undercomplete autoencoder, or the dimensions are increased, in what is called an overcomplete autoencoder. Hence, this idea is related to representation learning, where latent representations of the data are learned, either by obtaining a compressed version of the data or an overexpressed version. Variational graph autoencoders (VGAE) [110] appeared to perform edge-level prediction tasks by making use of graph convolution layers to produce the latent representation of the data. In fact, graph convolution layers are the building blocks of GAEs, however, other layers can be used [111]. Graph variational autoencoders (GVAE) [112] also make use of a GCN for the encoder. Graph autoencoders typically use a graph convolution approach to produce the latent features and the decoder part can use some well-known neural network layers to produce the output. An example is the marginalized graph autoencoder [113] that combines the use of graph convolutional layers with spectral clustering. Adversarially regularized graph autoencoders (ARGA) [114] make use of a graph convolution-based encoder and a graph embedding-based decoder to perform link prediction. Dirichlet graph variational autoencoders (DGVAE) [115] use Dirichlet priors and spectral clustering to enhance variational autoencoders. All in all, GAEs can be implemented by means of the layers seen so far, combining ideas from representation learning and dimensionality reduction to perform tasks at different levels, e.g., signal reconstruction, link prediction.
@@ 3.3.4 Recurrent GNNs
Recurrent GNNs usually combine GCNs with recurrent modules to take into account the temporal patterns of the data [116]. For instance, graph convolutional recurrent neural networks (GCRNs) put LSTM and the ChebNet together. Diffusion convolutional recurrent neural networks (DCRNNs) [117] combined diffusion graph convolutional layers with gated recurrent units. Gated graph neural networks (GGNNs) [118] used gated recurrent units with spatial convolutions to implement the recurrence [26]:
$\mathbf{h}^{(k)}(n)=\textbf{GRU}\left(\mathbf{h}^{(k-1)}(n),agg_{u\in\mathcal{N}(n)}\left(\mathbf{h}^{(k-1)}(u),\mathbf{W}\right)\right)$ (25)
where the hidden representation of node $n$ is a gated recurrent unit applied to the previous state of the $n$-th node ($\mathbf{h}^{k-1}(n)$) and a combination of the neighbors’ previous representations $agg_{u\in\mathcal{N}(n)}(\mathbf{h}^{(k-1)}(u),\mathbf{W})$. $\mathbf{W}$ are learnable parameters. These are just some examples of recurrent neural networks and how these are implemented by means of convolution operations together with classic recurrent units.
@@
Throughout this section we have reviewed the foundations of graph signal processing, machine learning over graphs and GNNs. In the next section, we review articles that apply many of the fundamentals we have seen to improve data quality in the context of monitoring sensor networks using graph models.
@@ 4 Graph-Based Data Quality Applications
Figure 10: Proposed taxonomy for the different data quality-enhancing tasks that can be carried out in a monitoring sensor network and the tools that can be used. The different tasks are classified according to the level to which they are applied; network, edge, and node level.
Figure 11: Different task levels for graph-based techniques in monitoring sensor networks. Network level tasks produce estimations for the network as a whole, edge level tasks produce estimations for the graph edges, and node level tasks produce estimations of graph signal values at different graph nodes.
As we have reviewed in the previous sections, many graph-based techniques allow the modeling of sensor network measurements [17]. There is a wide variety of applications where graphs can be leveraged to perform data quality-enhancing tasks, such as outlier detection or virtual sensing among others. Before delving into specific use cases, let us roughly classify the tasks according to their scope:
@@
Figure 11 represents the different task levels. Moreover, Figure 10 presents the data quality-enhancement task taxonomy described in this section, as well as the theoretical aspects used for each one of the applications - more details are given in the following sections. At the network level, we find tasks such as the detection of outliers and anomalies. At the edge level, we find tasks such as edge prediction and graph learning. At last, at the node level, we find applications such as outlier detection, missing value imputation, virtual sensing, or clustering. All these applications can be used to tackle data quality issues in order to produce continuous, complete, and accurate measurements [32].
@@ 4.1 Network Level
Let’s define a network level task as finding the map $f_{\mathcal{G}}{:\;}\mathbb{R}^{N\times T}{\rightarrow\;}\mathbb{R}$ such that $y{=}f_{\mathcal{G}}(\mathbf{X})$ where $\mathbf{X}{\in}\mathbb{R}^{N\times T}$ is a graph signal ($T{=}1$) or a set of graph signals ($T{>}1$). The target variable is $y$, which can be a categorical variable (e.g., classification or outlier detection task) or a scalar value depicting a metric for the graph.
4.1.1 Anomaly and Outlier Detection
@@
In the case of outlier and anomaly detection at the network level, we are interested in classifying a graph signal or a set of graph signals, i.e., $y{\in}\{1,\dots C\}$ where $C{\in}\mathbb{N}^{+}$ is the number of possible classes. For instance, the case of detecting whether the network measurements in a time instant correspond to an anomalous graph signal or not ($y{\in}\{0,1\}$). Hence, this approach does not provide localization capabilities, i.e., identifying the suspicious sensor or measurements that are causing the anomaly. Different machine learning approaches fit this setting; supervised machine learning, unsupervised learning, and semi-supervised learning. Supervised and semi-supervised models are limited in this field since labeled data about faulty sensors and abnormal measurements is difficult to obtain. The most common approach is the unsupervised learning setting where no labeled data is used and the data distribution together with the graph topology and a model are used. Ideally, data points, e.g., graph signals, that deviate from the distribution of the majority of observed graph signals are assumed to be anomalous.
@@
GSP metrics such as total variation or graph signal smoothness have been used to identify anomalous graph signals, e.g., $S(\mathbf{S},\mathbf{x}){>}TH$, where $TH{\in}\mathbb{R}$ is a threshold value indicating the graph signals presenting discrepancies between connected sensors which are prone to be anomalous [59, 61]. Other GSP methods use the magnitude of the GDFT high-frequencies, the spectral graph wavelet transform, or the GDFT representation of a graph filter output to detect anomalous signals, e.g., $\|\hat{\mathbf{x}}_{\text{high}}\|{>}TH$, [42, 121, 46, 60]. For instance, Xiao et al. [46] detected anomalies in a temperature monitoring IoT network by analyzing the frequency response of a graph filter. Figure 12 shows a real temperature monitoring network example, where the Laplacian matrix is used to obtain the smoothness and the GDFT, which can be later used to detect anomalous signals.
@@
Figure 12: Example of a temperature monitoring network in the area of Catalonia, Spain, with a total of 71 nodes. A KNN graph is constructed based on geographical distances. Afterward, it is shown how the signal smoothness $S(\mathbf{L},\mathbf{x})$ is increased due to a perturbation in a given graph signal, allowing the detection of an outlying signal. What is more, the norm of the GDFT coefficients associated with high frequencies, $\|\hat{\mathbf{x}}_{10:}\|$, can also be used to detect discrepancies in the measurements.
@@
Table 2: Examples of anomaly and outlier detection models based on graphs applied to network level tasks.
Reference Year Technique Use Case
Smoothness Chen et al. [61] 2015 Total Variation (TV) Temperature sensor network
Gopalakrishnan et al. [59] 2019 Total Variation (TV) Delays US airports
@@
Sandryhaila et al. [60] 2014 High-Frequencies GDFT Temperature sensor network
@@
GCN & GAE Chen et al. [45] 2021 GCN & Attention Mechanism Water sensor network
@@
Recently, Su et al. [21] have developed a graph-frequency domain filter which together with a Kalman filter is able to filter out anomalous graph signals in industrial pipe networks. Residual-based models have also been used to detect anomalies, where a graph signal reconstruction model, e.g., graph filter or GNN, is fitted with the data and large residuals’ norm depicts a possible outlier, $\|\mathbf{x}-\mathbf{f}_{\mathcal{G}}(\mathbf{x})\|{>}TH$. In this line, GNNs and GCNs have also been used as autoencoders to detect anomalous graph signals focusing on both signal reconstruction residuals and the reconstructed signal spectrum [125, 126]. Moreover, spatiotemporal GNNs have been used to detect anomaly signals in multivariate time series, including anomalies in water monitoring sensor networks [45]. Lin et al. [123] propose a context-aware graph autoencoder to detect anomalous air pollution events in an air quality monitoring network in Beijing. Feng et al. [122] propose a GAE based on graph convolutional layers to reconstruct the graph edges and graph signals and compute the anomaly score based on the loss function and the kernel density estimation of the distribution of scores during the training. Xu et al. [124] propose an edge-based GCN to detect anomalous electricity consumption records in smart meter networks in a graph classification supervised manner.
All in all, residual-based models are used to detect anomalies and outliers in an unsupervised manner, the signal reconstruction can be implemented in different ways, e.g., convex optimization, graph filters, GNN, and the anomaly score is computed based on some norm or distance. Graph filters can provide a scalable, interpretable, and convex alternative to highly complex GNNs. Nevertheless, depending on the data availability, GNNs can provide superior performance due to their ability to model multivariate time series, including temporal components, and complex surfaces. In any case, as we will see in section 5.1, GNNs and GSP tools can provide benefits in terms of model generalization and domain transferability, so easing the training of models and dealing with limited data scenarios. Table 2 lists some graph-based models as well as their use cases.

### v2

@@ 3.3.1 GNNs: Building Modules
3.3.2 Convolutional GNNs
@@
Figure 10: Example of a GCN architecture at the node level, where a graph signal $\mathbf{x}$ is reconstructed $\tilde{\mathbf{x}}$. The inputs are the graph signal and the graph shift matrix $\mathbf{S}$, and $\sigma(\cdot)$ denotes an activation function. A GNN architecture can include graph convolution and pooling layers, as well as readout layers and classic layers, e.g., fully-connected or convolutional layers.
@@
where $\mathbf{h}^{(k)}_{c_{k}}{\in}\mathbb{R}^{N}$ is the $c_{l}$ (channel) hidden state at layer $k$, $\sigma(\cdot)$ is the activation function, $c_{k}{\in}\mathbb{N}^{+}$ is the number of channels in layer $k$, $agg(\cdot)$ is an aggregation function used to aggregate the different filtered versions of graph signal $\mathbf{h}_{i}^{(k-1)}{\in}\mathbb{R}^{N}$ (e.g., sum or average), and $\boldsymbol{\theta}_{ij}^{(k)}{\in}\mathbb{R}^{K_{ij}}$ is the convolution function associated with a set of learnable parameters. Different types of networks and convolutions assume a different parametrization of the learnable parameters $\boldsymbol{\theta}$. For instance, the GCN [103] and Chebnet [39] use the Chebyshev polynomials to approximate the filtering, the CayleyNet [110] uses the Cayley polynomials to approximate the filtering, the Graph Wavelet Neural Network (GWNN) [111] uses wavelets to approximate the convolution operation, and the graph diffusion neural network (GDNN) [112] uses a graph diffusion convolution to implement the message passing operation, combining strengths of spatial and spectral methods. For instance, the implementation of the spectral graph convolution using the truncated expansion of the Chebyshev polynomials results in [39]:
@@
where $f(\cdot)$ is a function that combines the value at node $n$ with a function of the vicinity, $agg_{\forall i}(\cdot)$ is a function that aggregates the information of neighboring nodes of $\mathcal{N}(n)$, and ||||| | denotes a concatenation operation. Function $f$ combines the aggregated spatial information at node $n$ ($\mathbf{a}^{(k-1)}(n)$) with the previous hidden state $\mathbf{h}^{k-1}(n)$ at node $n$. Both $agg$ and $f$ can contain a set of learnable parameters $\boldsymbol{\theta}_{i}^{agg}$ and $\boldsymbol{\theta}_{j}^{f}$. This framework provides great generalization capabilities allowing the definition of aggregation and combination functions according to the data. The GAT [116] make use of the node spatial notion and graph structure to define the graph-based analogs of attention mechanisms.
3.3.3 Autoencoder GNNs
GAEs are widely used in applications such as outlier detection and missing value imputation. GAEs are trained to reconstruct the input signal, and in the process, a latent representation of the data is learned. To force the neural network to learn a latent representation, the dimensions of the hidden features of the encoder are decreased to produce a bottleneck effect, in what is called an undercomplete autoencoder, or the dimensions are increased, in what is called an overcomplete autoencoder. Hence, this idea is related to representation learning, where latent representations of the data are learned, either by obtaining a compressed version of the data or an overexpressed version. Variational graph autoencoders (VGAE) [117] appeared to perform edge-level prediction tasks by making use of graph convolution layers to produce the latent representation of the data. In fact, graph convolution layers are the building blocks of GAEs, however, other layers can be used [118]. Graph variational autoencoders (GVAE) [119] also make use of a GCN for the encoder. Graph autoencoders typically use a graph convolution approach to produce the latent features and the decoder part can use some well-known neural network layers to produce the output. An example is the marginalized graph autoencoder [120] that combines the use of graph convolutional layers with spectral clustering. Adversarially regularized graph autoencoders (ARGA) [121] make use of a graph convolution-based encoder and a graph embedding-based decoder to perform link prediction. Dirichlet graph variational autoencoders (DGVAE) [122] use Dirichlet priors and spectral clustering to enhance variational autoencoders. All in all, GAEs can be implemented by means of the layers seen so far, combining ideas from representation learning and dimensionality reduction to perform tasks at different levels, e.g., signal reconstruction, and link prediction.
@@ 3.3.4 Recurrent GNNs
Recurrent GNNs usually combine GCNs with recurrent modules to take into account the temporal patterns of the data [123]. For instance, graph convolutional recurrent neural networks (GCRNs) put LSTM and the ChebNet together. Diffusion convolutional recurrent neural networks (DCRNNs) [124] combine diffusion graph convolutional layers with gated recurrent units. Gated graph neural networks (GGNNs) [125] use gated recurrent units with spatial convolutions to implement the recurrence [30]:
$\mathbf{h}^{(k)}(n)=\textbf{GRU}\left(\mathbf{h}^{(k-1)}(n),agg_{u\in\mathcal{N}(n)}\left(\mathbf{h}^{(k-1)}(u),\boldsymbol{\theta}\right)\right)$ (22)
where the hidden representation of node $n$ is a gated recurrent unit applied to the previous state of the $n$-th node ($\mathbf{h}^{k-1}(n)$) and a combination of the neighbors’ previous representations $agg_{u\in\mathcal{N}(n)}(\mathbf{h}^{(k-1)}(u),\boldsymbol{\theta})$. $\boldsymbol{\theta}$ are learnable parameters. These are just some examples of recurrent neural networks and how these are implemented by means of convolution operations together with classic recurrent units.
@@
3.3.5 Generative GNNs
Generative models have attracted attention in recent years. Specifically, generative models can generate new synthetic measurements, or more complex structures such as graphs and sub-graphs. There are two main classes of models that are the most widely used, the variational autoencoders and the generative adversarial networks (GANs). In the context of graph variational autoencoders, models have been developed that make use of GCNs as encoders and learn a latent generative model $q(h)$ (where $h$ is the latent representation of the data) that allows for generating new instances of the graph shift matrix or graph signals [117, 128]. For example, Do et al. [129] use a VGAE consisting of an encoder formed by stacked layers of GCNs, a latent model parametrized as a Gaussian distribution, and a decoder composed of stacked GCNs to complete air quality measurements. Another successful example of generative models are GANs in which a sample generator $g$ competes with a discriminator whose goal is to distinguish real instances from generated instances in a zero-sum game [121, 130]. For example, VGAEs can be used as generators to generate synthetic samples or graphs [121] or more customized techniques can be developed to generate possible graph shift matrices [130].
Throughout this section, we have reviewed the foundations of GSP, ML over graphs and GNNs. In the next section, we review articles that apply many of the fundamentals we have seen to improve data quality in the context of monitoring sensor networks using graph models.
@@ 4 Graph-Based Data Quality Applications
Figure 11: Different task levels for graph-based techniques in monitoring sensor networks; network level tasks, edge level tasks, and node level tasks.
As we have reviewed in previous sections, many graph-based techniques allow the modeling of sensor network measurements [17]. There is a wide variety of applications where graphs can be leveraged to perform data quality-enhancing tasks, such as outlier detection or virtual sensing among others. Before delving into specific use cases, let us roughly classify the tasks according to their scope:
@@
Figure 12: Proposed taxonomy for the different data quality-enhancing applications that can be carried out in a monitoring sensor network and the techniques that can be used.
Figure 11 represents the different task levels. Moreover, Figure 12 presents the data quality-enhancement task taxonomy described in this section, as well as the theoretical aspects used for each one of the applications - more details are given in the following sections. At the network level, we find tasks such as the detection of outliers and anomalies which can be critical in assessing data consistency and coherence. At the edge level, we find tasks such as edge prediction and graph learning which might help in identifying correct graph structures and further applications to improve data accuracy. At last, at the node level, we find applications such as outlier detection to improve data consistency and coherence by identifying sensor measurement deviations, missing value imputation to ensure data completeness, virtual sensing to tackle different data quality dimensions such as accuracy or completeness, or clustering which can be useful in grouping the data and improving data accuracy in further application. All these applications can be used to tackle data quality issues to produce continuous, complete, and accurate measurements [36].
@@ 4.1 Network Level
Let’s define a network level task as finding the map $f_{\mathcal{G}}{:\;}\mathbb{R}^{N\times T}{\rightarrow\;}\mathbb{R}$ such that $y{=}f_{\mathcal{G}}(\mathbf{X})$ where $\mathbf{X}{\in}\mathbb{R}^{N\times T}$ is a graph signal ($T{=}1$) or a set of graph signals ($T{>}1$). The target variable is $y$, which can be a categorical variable (e.g., classification or outlier detection task) or a scalar value representing a metric for the graph.
4.1.1 Anomaly and Outlier Detection at Network Level
@@
In the case of outlier and anomaly detection at the network level, we are interested in classifying a graph signal or a set of graph signals, i.e., $y{\in}\{1,\dots C\}$ where $C{\in}\mathbb{N}^{+}$ is the number of possible classes. For instance, the case of detecting whether the network measurements in a time instant correspond to an anomalous graph signal or not ($y{\in}\{0,1\}$). Hence, this approach does not provide localization capabilities, i.e., identifying the suspicious sensor or measurements that are causing the anomaly. Different ML approaches fit this setting; supervised ML, unsupervised learning, and semi-supervised learning. Supervised and semi-supervised models are limited in this field since labeled data about faulty sensors and abnormal measurements is difficult to obtain. The most common approach is the unsupervised learning setting where no labeled data is used and the data distribution together with the graph topology and a model are used. Ideally, data points, e.g., graph signals, that deviate from the distribution of the majority of observed graph signals are assumed to be anomalous.
@@
GSP metrics such as total variation or graph signal smoothness have been used to identify anomalous graph signals, e.g., $S(\mathbf{S},\mathbf{x}){>}TH$, where $TH{\in}\mathbb{R}$ is a threshold value indicating the graph signals presenting discrepancies between connected sensors which are prone to be anomalous [65, 67]. Other GSP methods use the magnitude of the GDFT high-frequencies, the spectral graph wavelet transform, or the GDFT representation of a graph filter output to detect anomalous signals, e.g., $\|\hat{\mathbf{x}}_{\text{high}}\|{>}TH$, [46, 131, 50, 66]. For instance, Xiao et al. [50] detected anomalies in a temperature monitoring IoT network by analyzing the frequency response of a graph filter.
Recently, Su et al. [21] have developed a graph-frequency domain filter which together with a Kalman filter is able to filter out anomalous graph signals in industrial pipe networks. Residual-based models have also been used to detect anomalies, where a graph signal reconstruction model, e.g., graph filter or GNN, is fitted with the data and large residuals’ norm depicts a possible outlier, $\|\mathbf{x}-\mathbf{f}_{\mathcal{G}}(\mathbf{x})\|{>}TH$. Similarly, GNNs and GCNs have also been used as autoencoders to detect anomalous graph signals focusing on both signal reconstruction residuals and the reconstructed signal spectrum [132, 133]. Moreover, spatiotemporal GNNs have been used to detect anomaly signals in multivariate time series, including anomalies in water monitoring sensor networks [49]. Lin et al. [134] propose a context-aware graph autoencoder to detect anomalous air pollution events in an air quality monitoring network in Beijing. Feng et al. [135] propose a GAE based on graph convolutional layers to reconstruct the graph edges and graph signals and compute the anomaly score based on the loss function and the kernel density estimation of the distribution of scores during the training. Xu et al. [136] propose an edge-based GCN to detect anomalous electricity consumption records in smart meter networks in a graph classification supervised manner.
@@
Figure 13: Example of a temperature monitoring network in the area of Catalonia, Spain, with a total of 71 nodes. The signal smoothness, $S(\mathbf{L},\mathbf{x})$, and the norm of the GDFT coefficients associated to high-frequency components, $\|\hat{\mathbf{x}}_{10:}\|$, are used to detect anomalous measurements.
[…]
@@
Table 2: Examples of anomaly and outlier detection graph-based models applied at the network level.
Reference Year Method Use Case
GSP Smoothness Chen et al. [67] 2015 TV Temperature sensor network
Gopalakrishnan et al. [65] 2019 Delays US airports
@@
Sandryhaila et al. [66] 2014 High-Frequencies GDFT
@@
GNN GCN & GAE Chen et al. [49] 2021 GCN & Attention Mechanism Water sensor network
@@
All in all, residual-based models are used to detect anomalies and outliers in an unsupervised manner, the signal reconstruction can be implemented in different ways, e.g., convex optimization, graph filters, GNN, and the anomaly score is computed based on some norm or distance. Graph filters can provide a scalable, interpretable, and convex alternative to highly complex GNNs. Nevertheless, depending on the data availability, GNNs can provide superior performance due to their ability to model multivariate time series, including temporal components and complex surfaces. In any case, as we will see in section 5.2.1, GNNs and GSP tools can provide benefits in terms of model generalization and domain transferability, so easing the training of models and dealing with limited data scenarios. Table 2 lists some graph-based models as well as their use cases.